# Differential Privacy of Network Parameters from a System Identification Perspective

### v1

@@ 2 Problem Formulation
Additionally we need to define a notion of adjacency. We consider two GSO’s $\bm{S}$} and $\bm{S}$$^{\prime}$ to be adjacent if an only if
$\left\lVert{\bm{S}}-{\bm{S}}^{\prime}\right\rVert_{2}\leq\Delta_{{\bm{S}}}~~~\text{and}~~~\operatorname{rank}({\bm{S}}-{\bm{S}}^{\prime})\leq 1.$ (6)
Equation (6) requires that graphs only have a single edge change and that the spectrum of this change is bounded. In the traditional DP scenario, adjacent datasets are typically defined as having a single row change with bounded norm [12], and thus restricting a GSO to bounded single edge changes is the appropriate parallel [18, 19]. Finally, to align our scenario within the PDP framework, we let ${\bm{p}}(\operatorname{vec}\left({\bm{Y}}\right)|{\bm{S}};\bm{\phi})$ refer to the pdf of $\bm{Y}$} given GSO $\bm{S}$} and parameter vector $\bm{\phi}$ for $f_{{\bm{U}}}(\cdot;\bm{\phi})$. We then define
@@ 3 Analysis
$\Pr\left({\mathcal{L}}_{{\bm{S}},{\bm{S}}^{\prime}}(\cdot)>\epsilon\right)\leq\inf_{\alpha}\exp\left\{(\alpha-1)D_{\alpha}\left(\left.P\right|\left|Q\right.\right)-\alpha\epsilon\right\}.$ (10)
@@
$\inf_{\alpha}e^{-\alpha\epsilon}\mathbb{E}\left[e^{\alpha{\mathcal{L}}_{{\bm{S}},{\bm{S}}^{\prime}}(\cdot)}\right]=\inf_{\alpha>1}\exp\left\{(\alpha-1)D_{\alpha}\left(\left.P\right|\left|Q\right.\right)-\alpha\epsilon\right\}$
@@
$\dfrac{\alpha-1}{\alpha}D_{\alpha}+\dfrac{1}{\alpha}\ln\left(\frac{1}{\delta}\right)\leq\epsilon,$ (19)
@@ 3.1 Notes on Singular Matrices
$~~~~=\dfrac{{\bm{1}}_{\left\{\bar{{\bm{u}}}\in{\bm{W}}\right\}}}{\det{\bm{\Sigma}_{r}}}\int_{\mathbb{R}^{Tn-r}}f_{{\bm{U}}}\left({\bm{V}}_{r}\bm{\Sigma}_{r}^{-1}{\bm{Q}}_{r}^{\top}\bar{{\bm{y}}}+{\bm{V}}_{0}{\bm{w}};\theta\right)d{\bm{w}}.$

### v2

@@ 2 Problem Formulation
Additionally we need to define a notion of adjacency. We consider two GSO’s $\bm{S}$} and $\bm{S}$$^{\prime}$ to be adjacent if they differ by a single edge and
$\left\lVert{\bm{S}}-{\bm{S}}^{\prime}\right\rVert_{2}\leq\Delta_{{\bm{S}}}.$ (6)
In the traditional DP scenario, adjacent datasets are typically defined as having a single row change with bounded norm [12], and thus restricting a GSO to bounded single edge changes is the appropriate parallel [18, 19]. Finally, to align our scenario within the PDP framework, we let ${\bm{p}}(\operatorname{vec}\left({\bm{Y}}\right)|{\bm{S}};\bm{\phi})$ refer to the pdf of $\bm{Y}$} given GSO $\bm{S}$} and parameter vector $\bm{\phi}$ for $f_{{\bm{U}}}(\cdot;\bm{\phi})$. We then define
@@ 3 Analysis
$\Pr\left({\mathcal{L}}_{{\bm{S}},{\bm{S}}^{\prime}}(\cdot)>\epsilon\right)\leq\inf_{\alpha}\exp\left\{(\alpha-1)\left(D_{\alpha}\left(\left.P\right|\left|Q\right.\right)-\epsilon\right)\right\}.$ (10)
@@
$\inf_{\alpha}e^{-\alpha\epsilon}\mathbb{E}\left[e^{\alpha{\mathcal{L}}_{{\bm{S}},{\bm{S}}^{\prime}}(\cdot)}\right]=\inf_{\alpha>1}\exp\left\{(\alpha-1)\left(D_{\alpha}\left(\left.P\right|\left|Q\right.\right)-\epsilon\right)\right\}$
@@
$D_{\alpha}+\dfrac{1}{\alpha-1}\ln\left(\frac{1}{\delta}\right)\leq\epsilon,$ (19)
@@ 3.1 Notes on Singular Matrices
$~~~~=\dfrac{{\bm{1}}_{\left\{\bar{{\bm{y}}}\in{\bm{W}}\right\}}}{\det{\bm{\Sigma}_{r}}}\int_{\mathbb{R}^{Tn-r}}f_{{\bm{U}}}\left({\bm{V}}_{r}\bm{\Sigma}_{r}^{-1}{\bm{Q}}_{r}^{\top}\bar{{\bm{y}}}+{\bm{V}}_{0}{\bm{w}};\theta\right)d{\bm{w}}.$